OpenStudy (anonymous):

What is the product of five over six and 2? one and two over three one and one over three five over twelve ten over twelve

OpenStudy (ashleyisakitty):

Do you mean like 5/6+2?

OpenStudy (anonymous):

yes

OpenStudy (ashleyisakitty):

5/6+2=17/6, and as a mixed fraction its 2 and 5/6. I dont see this in any of your choices... sorry :/

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

i think its 5/6 times 2 because you said product and that means multiplication so (5/6)(2/1)= 10/6= one and 2/3

OpenStudy (anonymous):

@messiahcollins

OpenStudy (anonymous):

I see what your doing

OpenStudy (anonymous):

okay good just making sure